SPECTROSCOPY OF ACETYLENE TRIPLET STATES BY DIRECT OPTICAL EXCITATION

\underline{R. L. THOM} and R. W. FIELD, Department of Chemistry, Massachusetts Institute of Technology, Cambridge, MA 02139.

We present here initial results of the first experiments detecting direct optical (uv) excitation of spin-forbidden transitions from the singlet ground electronic surface (S_{0}) of acetylene into the low-lying molecular triplet manifold (T_{1},T_{2}). The metastable states thus populated are detected via electron ejection from a Cs metal surface (SEELEM: Surface Electron Ejection by Laser Excited Metastables spectroscopy). Initial insights into the structural and dynamical properties of these states, and the mechanism of the SEELEM detection process, are discussed.